OpenStudy (calculusxy):

MEDAL WILL BE AWARDED!!! PLEASE HELP Find the mean absolute deviation for the following set of high temperatures recorded across the United States on August 21. {87, 72, 93, 91, 64, 86, 95}

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What is the mean of this data set?

OpenStudy (calculusxy):

I got the mean of 84.

OpenStudy (calculusxy):

I seemed to do that step well, but then when it came to adding all the absolute values, the mean keeps on ending up as a decimal.

jimthompson5910 (jim_thompson5910):

now subtract each value from the mean make those differences positive (so use absolute value)

OpenStudy (calculusxy):

okay, so maybe you may notice the mistake, if i did get one. 87-84=3 72-84=-8 93-84=9 91-84=7 64-84=-20 86-84=2 95-84=11 so i added the absolute values: 3+8+9+7+20+2+11=60 i am pretty sure that i got that one incorrect. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

72-84 should be -12 not -8

OpenStudy (calculusxy):

the answer is 12 right?

jimthompson5910 (jim_thompson5910):

that's not the final answer

jimthompson5910 (jim_thompson5910):

I'm just correcting part of your scratch work

OpenStudy (calculusxy):

wait, i am kind of confused in here. am i correct?

jimthompson5910 (jim_thompson5910):

go back in how you subtracted in your work, you wrote "72-84=-8" when you *should* have this "72-84=-12"

OpenStudy (calculusxy):

yes that's what i did. then i added all the numbers up, and then divided it by 7 to get the mean, and got 12.

jimthompson5910 (jim_thompson5910):

the answer is still not 12

OpenStudy (calculusxy):

so what's the next step?

jimthompson5910 (jim_thompson5910):

you add up the absolute value differences (ie the distance each value is from the mean) then divide by 7

jimthompson5910 (jim_thompson5910):

here's a page with an example http://www.mathsisfun.com/data/mean-deviation.html

OpenStudy (calculusxy):

i did and got the answer as 84, divided it by 7 = 12?

jimthompson5910 (jim_thompson5910):

look over that page I posted

OpenStudy (calculusxy):

i am not getting this. so please can you tell me where i got my error?

OpenStudy (calculusxy):

please hurry up i have to go somewhere. ..

jimthompson5910 (jim_thompson5910):

you have these distances: 3,12,9,7,20,2,11 average them to find your answer

OpenStudy (calculusxy):

so i add all of them and then divide it by 7?

OpenStudy (calculusxy):

please hurry ...

jimthompson5910 (jim_thompson5910):

yes, add them up and divide by 7